Help would be much appreciated. The length of the side opposite of the 30˚of a 30˚-60˚-90˚ triangle is 12.7 centimeters. Find the length of the hypotenuse. (Answer with just the number- no label)
Answer by rothauserc(4718)

let x be the hypotenuse
sin(30) = 12.7 / x
x = 12.7 / sin(30) = 12.7 / (1/2) = 25.4